OpenStudy (anonymous):

Please can someone HELP me on this. Match the following items by evaluating the expression for x = -2. 1. x^-2 ___ -1/2 2. x^-1 ___ 4 3. x^0 ___ -2 4. x^1 ___ 1/4 5. x^2 ___ 1

OpenStudy (anonymous):

\[\LARGE x^{-2}=\frac{1}{x^2}=\frac{1}{(-2)^2}=?\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

how about this: \[\LARGE x^1=(-2)^1=?\]

OpenStudy (anonymous):

i kind of asked to match them ...

OpenStudy (anonymous):

yes but you have to solve them before matching...

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

... first one is... \[\LARGE x^{-2}=\frac{1}{x^2}=\frac{1}{(-2)^2}=\frac14 \]

OpenStudy (anonymous):

so the last one \[\LARGE x^2=(-2)^2=?\]